Calculate the weighted average outstanding shares. The answer is 249.

 

1/1 Outstanding = 100 shares

4/1 Issued = 20 shares

6/1 Declared dividend = 10%

10/1 Acquired = 10 shares

11/1 declared = 2:1 stock split


1
Expert's answer
2022-02-28T11:43:59-0500

Outstanding shares

No of shares. Weight No

"A=100\times1=100"


"B=(100+20)\times0.75=90"


"C=120\times0.25=30"


"D=165\times0.17=28"


"Total=A+B+C+D"


"Total=100+90+30+28=248"


Total weighted shares are 248.